OpenStudy (anonymous):

Find the perimeter of a triangle with vertices (2,4) (-1,3) and (2, -3). (show your work on graph paper) and solve 3^4x-7=4^2x+3

OpenStudy (anonymous):

perimer = distance problem

OpenStudy (anonymous):

let A be (2,4), B (-1,3), C (2,-3) the three vertices of your triangle. \[ \begin{align} & |AB| = \sqrt{(3-4)^2+(-1-2)^2} \\ &|BC| = \sqrt{(-3-3)^2+(2-(-1))^2} \\ & |AC| = \sqrt{(4-(-3))^2+(2-2)^2} \end{align}\] do the computation and add the three lenght to get the perimeter. the general formula for two point \(A: (x_a,y_a)\ and\ B:(x_b,y_b) \) is: \[ |AB| = \sqrt{(y_b-y_a)^2+(x_b-x_a)^2}\]

OpenStudy (anonymous):

The second part of your exo has been answered allready here : http://openstudy.com/updates/4f402100e4b0534c53cc0016

OpenStudy (anonymous):

@mkone shouldn't you use log here. what's the rules for logs again

OpenStudy (anonymous):

please check the link

OpenStudy (anonymous):

i checked the link but i dont understand why you can't use logarithm here. why did you use the natural log

OpenStudy (anonymous):

You can replace \(ln \ by\ log\) it will not change the final result nor the step of the demonstration. \(ln\) is just short for me to write